OpenStudy (anonymous):

integrate x tanh^-1 x dx

OpenStudy (anonymous):

First thing that comes to mind:\[ \tanh(x) = \frac{e^x-e^{-x}}{e^x+e^{-x}} \]So maybe if you solve for \(y\)\[ y = \frac{e^x-e^{-x}}{e^x+e^{-x}} \]You will get a function that isn't too hard to integrate.

OpenStudy (anonymous):

I mean if you solve for \(x\) .

OpenStudy (anonymous):

@wio what in my mind is only by parts or tabula maybe?

OpenStudy (anonymous):

If you did do it by parts, then you want \(dv\) to be easy to integrate. So you'd have to do something like \(u = tan^{-1}(x) \quad dv = xdx\).

OpenStudy (anonymous):

@wio tanh^-1 h there

OpenStudy (anonymous):

Yeah, oops.

OpenStudy (anonymous):

but, when i tried that way it seem my working does not stop

OpenStudy (anonymous):

Can you plug it into the formula then? \[ \int udv = uv-\int vdu \]

OpenStudy (anonymous):

Okay, what did you get so far?

OpenStudy (anonymous):

Your work, should definitely stop at some point.

OpenStudy (anonymous):

\[ u = \tanh^{-1}(x)\\ du = ? \\ v = ? \\ dv = xdx \]

OpenStudy (anonymous):

\[\frac{ x^2 }{ 2} \tanh^-1 - \frac{ x^2 }{ 2(1-2x) }\]

OpenStudy (anonymous):

i left integrate symbol there

OpenStudy (anonymous):

@wio \[\int\limits_{}^{}\frac{ x^2 }{ 2-2x62 }\]

OpenStudy (anonymous):

it should be x^2 there.,wat would u get?

OpenStudy (anonymous):

I'm getting \[ \int \frac{x^2dx}{2-2x^2} \]as my \(\int vdu\)

OpenStudy (anonymous):

What about doing another u substitution now? What about \(u = 1-x^2\)

OpenStudy (anonymous):

Or maybe just \(u^2 = 1-x^2\).

OpenStudy (anonymous):

there still x above

OpenStudy (anonymous):

Then \(x^2 = u^2+1\)?

OpenStudy (anonymous):

That would get you: \[ \int \frac{u^2+1}{2u^2}du = \int \frac{u^2}{2u^2} + \frac{1}{2u^2}du \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you see how to do it now?

OpenStudy (anonymous):

@gee can you do the rest?

OpenStudy (anonymous):

@wio i think i just want to use the substitution of identity

OpenStudy (kainui):

I learned DETAIL for an easy way to do integration by parts. Whatever letter is closer to D in detail should be the one you choose as the one multiplied by dx. I'll show you. D E - exponentials like e^x T - trig functions like sin(x) A - algebraic like x^2 I - inverse trig functions like tan^-1(x) L- logarithms like ln(x) So if you see integral of x*e^xdx then you would separate them out as: x=u and e^xdx=dv since e^x is closer to D in DETAIL.